How many liters of pure acetic acid must be mixed with 10 liters of a 30% acetic acid soluti0n to get a 60% soluti0n?
Answer by ptaylor(2198)

Let x=amount of pure acetic acid needed
Now we know that the amount of pure acetic acid that exists before the mixture takes place (x+0.30*10) has to equal the amount of pure acetic acid after the mixture takes place (0.60(x+10)). Sooooo
x+0.30*10=0.60(x+10) get rid of parens and simplify
x+3=0.60x+6 subtract 3 and also 0.60x from each side
x+3-3-0.60x=0.60x-0.60x+6-3 collect like terms
0.40x=3
x=7.5 liters----------------amount of pure acetic acid needed
CK
7.5+0.30*10=0.60*17.5
7.5+3=10.5
10.5=10.5
Hope this helps---ptaylor